How do you solve y^(-1/3) -3y^(2/3) /y^(-4/3)? No decimals or negative exponents please
Answer by stanbon(48569)

y^(-1/3) -3y^(2/3) /y^(-4/3)
Factor the numerator to get:
= y^(-1/3)[1 - 3y]/y^(-4/3)
Simplify to get:
= y(1-3y)
Change form:
= y - 3y^2
==============
Cheers,
Stan H.